3(4m-1)^2+7(4m-1)=-2
Answer by lwsshak3(11628)

3(4m-1)^2+7(4m-1)=-2
let u=(4m-1)
u^2=(4m-1)^2
..
3u^2+7u=-2
3u^2+7u+2=0
(3u+1)(u+2)=0
3u+1=0
u=-1/3=4m-1
4m=-1/3+1=2/3
m=2/12=1/6
..
u+2=0
u=-2=4m-1
4m=-1
m=-1/4
ans:
m=1/6 or -1/4